OpenStudy (anonymous):

Please help!! i will medal + fan Which equation shows the distributive property? A. -5 + (7+3) = (-5+7) + 3 B. -3 * (2*7) = (-3*2) * 7 C. -3 + (7+3) = -3 + (3+7) D. -5

OpenStudy (cocoacat):

B! For multiple choice, it often helps to get rid of answers that don't make sense immediately. For example, D doesn't show anything. C is showing a property of addition that order doesn't matter, and so does A. B is the only one that shows multiplication into a parentheses, which is the distributive property!

OpenStudy (welshfella):

a(b + c) = a*b + a*c is the distributive property

OpenStudy (welshfella):

also a(bc)= ab*ac

OpenStudy (welshfella):

I think there's a typo in one of the choices

OpenStudy (anonymous):

oh im sorry number D is actually " -5(2+9) = -5 * 2 + (-5) * 9 "

OpenStudy (anonymous):

@cocoacat @welshfella

OpenStudy (cocoacat):

Ah then it is number D, as the distributive property is that a(b+c) = ab + ac!